(1) Use inductive reasoning to predict the next number in each list.

(a) 1, 4, 9, 16, 25, 36, 49, ?

(b) 2, 7, −3, 2, −8, −3, −13, −8, −18, ?

(c) 1/2, 2/3, 3/4, 4/5, 5/6, 6/7, ?


1
Expert's answer
2021-10-17T16:49:40-0400

(1) Let us use inductive reasoning to predict the next number in each list.


(a) It follows that the sequence is defined as "x_n=n^2," and hence the next number is "8^2=64."


(b) It follows that "x_{2n}=5+x_{2n-1}" and "x_{2n+1}=x_{2n}-10," and hence the next number is "-13."


(c) It follows that the sequence is defined as "x_n=\frac{n}{n+1}," and hence the next number is "\frac{7}{8}."